# Unified Parameter-Efficient Unlearning for LLMs

### v1

@@ Abstract
The advent of Large Language Models (LLMs) has revolutionized natural language processing, enabling advanced understanding and reasoning capabilities across a variety of tasks. Fine-tuning these models for specific domains, particularly through Parameter-Efficient Fine-Tuning (PEFT) strategies like LoRA, has become a prevalent practice due to its efficiency. However, this raises significant privacy and security concerns, as models may inadvertently retain and disseminate sensitive or undesirable information. To address these issues, we introduce a novel instance-wise unlearning framework, LLMEraser, which systematically categorizes unlearning tasks and applies precise parameter adjustments using influence functions. Unlike traditional unlearning techniques that are often limited in scope and require extensive retraining, LLMEraser is designed to handle a broad spectrum of unlearning tasks without compromising model performance. Extensive experiments on benchmark datasets demonstrate that LLMEraser excels in efficiently managing various unlearning scenarios while maintaining the overall integrity and efficacy of the models.
@@ 1 Introduction
Prior efforts in exploring unlearning in LLMs primarily focus on removing specific concepts (Kassem et al., 2023; Jang et al., 2023). A typical example is the erasure of LLM’s ability to recall information related to the Harry Potter series (Eldan & Russinovich, 2023). While these efforts yield valuable insights, they risk inadvertently affecting related concepts, such as other novels with similar titles. In this work, we broaden the scope by investigating instance-wise unlearning tasks, which allow us to target more nuanced aspects of model behavior. To this end, we first present various instance-wise unlearning tasks for LLMs, as illustrated in Figure 1. More case studies can be found in Appendix B. Specifically, consider a training instance $z=(x,y)$ in a supervised fine-tuning dataset, where $x$ represents the query and $y$ is the response. We can categorize the LLMs unlearning tasks at the instance level as follows:
@@
In this work, we focus on unlearning the domain-specific data used solely in PEFT, which requires updating the PEFT adapters (e.g., LoRA). Technically, recent LLM-unlearning efforts can be roughly grouped into two categories. Exact unlearning approaches divide data into disjoint shards and retrain adapters (Bourtoule et al., 2021; Hu et al., 2024b). Despite effectiveness, these methods have inherent limitations — inevitably destroying the model’s original structure and necessitating the retraining cost. Approximate unlearning methods, on the other hand, aim to replicate the performance of the retrained model, often aligning the output of the target data closely with randomness through KL-divergence-based PEFT (Liu et al., 2024a; Qu et al., 2024). Nonetheless, this paradigm primarily focuses on data removal (e.g., IR) and hardly corrects biased or inaccurate data (e.g., QM, RC), as it falls short in guiding the output of the target data towards accurate information, rather than mere randomness. See Table 1 for the summary of current LLMs unlearning methods, with detailed descriptions available in Appendix A. Overall, both approaches struggle to efficiently handle these instance-wise LLM unlearning tasks and are not specifically designed for unlearning within the PEFT framework. It calls for a general LLM unlearning method capable of addressing these various tasks.
Table 1: A summary of existing LLM unlearning methods and their application scenarios. E and A are abbreviations for Exact unlearning and Approximate unlearning, respectively.
@@
In pursuit of parameter-efficient unlearning, we identify the influence function (Koh & Liang, 2017) as a promising tool. At its core is to formulate the parameter changes caused by perturbations in the form of the inverse-Hessian-vector product (Agarwal et al., 2016), where Hessian matrix represents the curvature of the loss function w.r.t. model parameters. However, the direct application of the influence function to LLMs presents two significant challenges: the expensive cost of calculating the inverse Hessian-vector product for vast model parameters and the cumulative errors introduced by approximation strategies (e.g., stochastic estimation (Agarwal et al., 2016)). Consequently, the use of influence functions for LLM unlearning remains largely underexplored. To fill this research gap, we propose a unified parameter-efficient unlearning framework, LLMEraser, for various instance-wise unlearning tasks. Specifically, for each type of unlearning task, LLMEraser leverages influence functions to directly calculate the parameter changes in the PEFT adapters and then efficiently update the adapter parameters, thus bypassing the need for time-consuming model retraining or fine-tuning. Furthermore, we reformulate the calculation of the inverse-Hessian-vector product into a finite-sum quadratic programming problem (Nesterov, 2013; Beck & Teboulle, 2009), significantly reducing computational complexity while mitigating the approximation errors from stochastic estimation. LLMEraser has several advantages: model-agnostic, applicable to various instance-wise unlearning tasks, and ensuring fast model updates. We conduct experiments on both LLMs and Multimodal Large Language Models (MLLMs), specifically focusing on LLMs for Recommendation (LLM4Rec) as well as MLLM relation mining tasks to validate the effectiveness of LLMEraser. Our extensive evaluations across these diverse scenarios demonstrate that LLMEraser consistently outperforms the state-of-the-art unlearning methods.
@@ 3.3 A New Algorithm for Computing Parameter Changes
LLMEraser reformulates the calculation of parameter changes as solving an equivalent optimization problem expressed in summation form, enabling efficient resolution using mini-batch algorithms. Specifically, we focus on the following optimization problem regarding $\Delta$:
@@
$f((x,y),\Delta)=\frac{1}{2}\Delta^{\top}\nabla_{\Theta}^{2}\mathcal{L}\left((x,y),\hat{\Theta}\right)\Delta+\langle b,\Delta\rangle.$ (15)
By employing scalable algorithms (e.g., SGD) to optimize problem 12, we can obtain the solution for $\Delta\Theta_{\text{Task}}$. It is worth noting that both the function value and the gradient can be efficiently computed using the Hessian-vector-product (HVP) [^1] HVP has a corresponding implementation in PyTorch; refer to https://pytorch.org/docs/stable/autograd.html for details., reducing the complexity from $\mathcal{O}(p^{2})$ to $\mathcal{O}(p)$ (Pearlmutter, 1994), where $p$ is the number of trainable parameters. The pseudocode for computing parameter changes can be found in Appendix C. Error analysis for our proposed algorithm can be found in Appendix E.
@@ 4.1 Experimental Setups
We conduct experiments on both LLMs and Multimodal Large Language Models (MLLMs), focusing specifically on LLMs for Recommendation (LLM4Rec) (Bao et al., 2023; Liao et al., 2024) and MLLM relation mining tasks (Wu et al., 2024; Ye et al., 2024), to validate the effectiveness of our proposed LLMEraser. We choose LLaMA2-7B (Touvron et al., 2023b) as our backbone LLM and LLaVA 1.5-7B (Liu et al., 2023a) for the MLLM experiments. Comprehensive details on task, datasets, baselines, and evaluation metrics for our proposed LLMEraser can be found in Appendix D.1.
@@ 4.4 Results analysis for Unlearning Efficiency (RQ3)
Efficiency is a key metric in evaluating unlearning techniques, particularly for LLMs. We here conduct experiments, comparing our proposed LLMEraser against existing techniques. For a fair comparison, we report the execution time in the QM task, where $5\%$ of users have items replaced with noisy interactions. All methods are run on a single Nvidia A100 GPU. Table 6 presents the results. We can observe that: (1) Due to the parallel training of sub-models, the retraining time of both SISA and RecEraser can be reduced to some extent. However, RecEraser requires data partitioning based on similarity, which introduces additional computational overhead. Moreover, both methods remain highly inefficient as unlearning requests necessitate retraining of the adapters. (2) In contrast, our proposed LLMEraser exhibits remarkable efficiency in handling unlearning tasks. By directly modifying model parameters, LLMEraser achieves a speedup of approximately 31.25 times compared to retraining, requiring only about $1.4\times 10^{3}$seconds to update the parameters. This reduction in execution time demonstrates the effectiveness of our approach in accelerating the computation of parameter changes. Additional experimental results and related analyses on the memory usage and execution time of LLMEraser can be found in Appendix F.
@@ 6 Conclusion And Future Work
Ethics Statement
This work is primarily foundational in instance-wise unlearning for LLMs, focusing on the development of a more efficient parameter-efficient unlearning framework. Its primary aim is to contribute to the academic community by enhancing the understanding and implementation of LLM applications in specific domains. We do not foresee any direct, immediate, or negative societal impacts stemming from the outcomes of our research.
Reproducibility Statement
All results presented in this work are fully reproducible. We provide the code for our method and baseline models in the supplementary materials. The optimal hyperparameters utilized in our experiments are detailed in Appendix D.2.
Appendix A Overview of existing LLM unlearning
@@
  • APA (Hu et al., 2024b): This study introduces the first exact unlearning approach for large language model-based recommendation (LLMRec), focusing on the removal of personal data to comply with privacy regulations. The Adapter Partition and Aggregation (APA) method is proposed, which combines data partitioning with parameter aggregation to reduce inference latency while maintaining performance. This approach enables efficient unlearning without incurring the extra costs typically associated with traditional methods. However,it can affect the integrity of the adapter structure and necessitates retraining of sub-models.
@@
Appendix B More Examples of various unlearning tasks
Figure 5: Response Correction Case Study.
Appendix C Algorithm for calculating parameter changes
@@
Appendix D Experimental Details
D.1 The tasks for LLMs and MLLMs
@@
  • In terms of MLLMs unlearning tasks, we focus on hard hallucination mining, e.g., understanding of relation (Wu et al., 2024) and spurious biases (Ye et al., 2024). We structure the evaluation as binary or multi-choice classification problems, which aim to select the ground-truth from the noisy labels. Specifically, for relation understanding, we follow (Wu et al., 2024) to present the Recall, F1-Score, Precision, Classification accuracy, Yes ratio as the evaluation metrics. For spurious biases, we follow (Wu et al., 2024) to show the classification accuracy for 9 types of spurious correlations, which is Background (BG), Texture and Noise (TN), Co-occurring Objects (CO), Relative Size (RS), Colorization (Col.), Orientation (Ori.), Lighting and Shadows (LS), Perspective and Angle (PA), and Shape (Sha.).
Datasets: Our experimental datasets for LLM4Rec unlearning tasks include three commonly used recommendation datasets: BookCrossing (Ziegler et al., 2005), MovieLens (Harper & Konstan, 2016), and LastFM (Cantador et al., 2011). We follow the data preprocessing and dataset partitioning as described in (Bao et al., 2023) and (Liao et al., 2024). For MLLMs unlearning tasks, we utilize MMSpuBench (Ye et al., 2024), and R-Bench (Wu et al., 2024) with the representative masked instances for evaluation, partitioning the data is into training (60%), validation (20%), and testing (20%) set.
@@
Table 7: The hyperparameter settings and details for PEFT on each dataset or task.
Method / Dataset LoRA r
LoRA
alpha
LoRA
dropout
Target modules #Params
#Trainable
params
Percentage of
trainable params
TALLRec 8 16 0.05 ’q proj’,’v proj’ $6.7\times 10^{10}$ $4.2\times 10^{7}$ 0.06%
LLaRA 8 16 0.10
’k proj’, ’v proj’,
’q proj’, ’o proj’,
’gate proj’,
’up proj’,
’down proj’
$6.7\times 10^{10}$ $1.9\times 10^{8}$ 0.30%
MM-SPUBENCH 8 16 0.05
’k proj’, ’v proj’,
’q proj’, ’o proj’,
’gate proj’,
’up proj’,
’down proj’
$7.1\times 10^{10}$ $4.1\times 10^{8}$ 0.58%
R-BENCH 8 16 0.05
’k proj’, ’v proj’,
’q proj’, ’o proj’,
’gate proj’,
’up proj’,
’down proj’
$7.1\times 10^{10}$ $4.1\times 10^{8}$ 0.58%
D.2 Implementation details
All implementations are performed using Python $3.8$ and PyTorch Lightning $1.8.6$. All methods are run on a single Nvidia A100 GPU.
For R-Bench, the learning rate is set to $41e-$ 4, with a batch size of $16$. The MLP model uses a learning rate of $52e-$ 5. During the unlearning phase, the parameters are specified as: $\Delta_{lr}=1e-5$ and $\Delta_{init}=1e-5$. In MMSPubench, the unlearning parameters are specified as: $\Delta_{lr}=2e-5$ and $\Delta_{init}=1e-5$.
For the TallRec model, we selected optimal training parameters based on the original work. During the unlearning phase, the parameters are configured as: $\Delta_{lr}=5e-5$ and $\Delta_{init}=5e-3$.
For the LLARA model, we selected optimal training parameters based on the original work. For the MovieLens dataset, the parameters are set as follows: $\Delta_{lr}$ is $41e-$ 4 and $\Delta_{init}$ is $51e-$ 5. For the LastFM dataset, the parameters are configured to $\Delta_{lr}$ at $51e-$ 5 and $\Delta_{init}$ at $61e-$ 6.
The hyperparameter settings and details for PEFT on each dataset or task can be found in Table 7.
Appendix E Estimation errors analysis of LLMEraser
The approximation errors in LLMEraser consist of two primary components: first, the errors introduced by the Taylor expansion approximation in the derivation of the influence function, where high-order terms are neglected; and second, the errors arising from the new algorithm proposed in Section 3.3 in our manuscript for solving the inverse Hessian-vector product. We will conduct the error analysis in two parts accordingly.
E.1 Errors analysis for Taylor expansion approximation
Without loss of generality, we consider approximation error in Equation 6 of our submitted manuscript. In other words, we will analyze the error $\|\Delta\Theta(\epsilon)+\epsilon H_{\hat{\Theta}}^{-1}\nabla_{\Theta}\mathcal{L}((x,y);\hat{\Theta})\|.$
The derivation below follows from (Zhang et al., 2022), where we assume that $H_{\hat{\Theta}}$ is invertiable. As we discussed in our manuscript, this can be guaranteed if the second-order sufficient condition holds at $\hat{\Theta}$.
Since $\hat{\Theta}_{\text{new}}(\epsilon)$ is an optimal solution to the perturbed loss function defined in Equation 5 in the submitted manuscript, we have

### v2

@@ Abstract
The advent of Large Language Models (LLMs) has revolutionized natural language processing, enabling advanced understanding and reasoning capabilities across a variety of tasks. Fine-tuning these models for specific domains, particularly through Parameter-Efficient Fine-Tuning (PEFT) strategies like LoRA, has become a prevalent practice due to its efficiency. However, this raises significant privacy and security concerns, as models may inadvertently retain and disseminate sensitive or undesirable information. To address these issues, we introduce a novel instance-wise unlearning framework, LLMEraser, which systematically categorizes unlearning tasks and applies precise parameter adjustments using influence functions. Unlike traditional unlearning techniques that are often limited in scope and require extensive retraining, LLMEraser is designed to handle a broad spectrum of unlearning tasks without compromising model performance. Extensive experiments on benchmark datasets demonstrate that LLMEraser excels in efficiently managing various unlearning scenarios while maintaining the overall integrity and efficacy of the models. Our code is available at https://github.com/oceanoceanna/LLMEraser.
@@ 1 Introduction
Prior efforts in exploring unlearning in LLMs primarily focus on removing specific concepts (Kassem et al., 2023; Jang et al., 2023). A typical example is the erasure of LLM’s ability to recall information related to the Harry Potter series (Eldan & Russinovich, 2023). While these efforts yield valuable insights, they risk inadvertently affecting related concepts, such as other novels with similar titles. In this work, we broaden the scope by investigating instance-wise unlearning tasks, which allow us to target more nuanced aspects of model behavior. To this end, we first present various instance-wise unlearning tasks for LLMs, as illustrated in Figure 1. More case studies can be found in Appendix G. Specifically, consider a training instance $z=(x,y)$ in a supervised fine-tuning dataset, where $x$ represents the query and $y$ is the response. We can categorize the LLMs unlearning tasks at the instance level as follows:
@@
In this work, we focus on unlearning the domain-specific data used solely in PEFT, which requires updating the PEFT adapters (e.g., LoRA). Technically, recent LLM-unlearning efforts can be roughly grouped into two categories. Exact unlearning approaches divide data into disjoint shards and retrain adapters (Bourtoule et al., 2021; Hu et al., 2024c). Despite effectiveness, these methods have inherent limitations — inevitably destroying the model’s original structure and necessitating the retraining cost. Approximate unlearning methods, on the other hand, aim to replicate the performance of the retrained model, often aligning the output of the target data closely with randomness through KL-divergence-based PEFT (Liu et al., 2024a; Qu et al., 2024). Nonetheless, this paradigm primarily focuses on data removal (e.g., IR) and hardly corrects biased or inaccurate data (e.g., QM, RC), as it falls short in guiding the output of the target data towards accurate information, rather than mere randomness. See Table 1 for the summary of current LLMs unlearning methods, with detailed descriptions available in Appendix A. Overall, both approaches struggle to efficiently handle these instance-wise LLM unlearning tasks and are not specifically designed for unlearning within the PEFT framework. It calls for a general LLM unlearning method capable of addressing these various tasks.
Table 1: A summary of existing LLM unlearning methods and their application scenarios. ‘E’ and ‘A’ are abbreviations for Exact unlearning and Approximate unlearning, respectively.
@@
In pursuit of parameter-efficient unlearning, we identify the influence function (Koh & Liang, 2017) as a promising tool. At its core is to formulate the parameter changes caused by perturbations in the form of the inverse Hessian-vector-product (Agarwal et al., 2016), where Hessian matrix represents the curvature of the loss function w.r.t. model parameters. However, the direct application of the influence function to LLMs presents two significant challenges: the expensive cost of calculating the inverse Hessian-vector-product for vast model parameters and the cumulative errors introduced by approximation strategies (e.g., stochastic estimation (Agarwal et al., 2016)). Consequently, the use of influence functions for LLM unlearning remains largely underexplored. To fill this research gap, we propose a unified parameter-efficient unlearning framework, LLMEraser, for various instance-wise unlearning tasks. Specifically, for each type of unlearning task, LLMEraser leverages influence functions to directly calculate the parameter changes in the PEFT adapters and then efficiently update the adapter parameters, thus bypassing the need for time-consuming model retraining or fine-tuning. Furthermore, we reformulate the calculation of the inverse Hessian-vector-product into a finite-sum quadratic programming problem (Nesterov, 2013; Beck & Teboulle, 2009), significantly reducing computational complexity while mitigating the approximation errors from stochastic estimation. LLMEraser has several advantages: model-agnostic, applicable to various instance-wise unlearning tasks, and ensuring fast model updates. We conduct experiments on both LLMs and Multimodal Large Language Models (MLLMs), specifically focusing on LLMs for Recommendation (LLM4Rec) as well as MLLM relation mining tasks to validate the effectiveness of LLMEraser. Our extensive evaluations across these diverse scenarios demonstrate that LLMEraser consistently outperforms the state-of-the-art unlearning methods.
@@ 3.3 A New Algorithm for Computing Parameter Changes
Inspired by the previous studies (Ding et al., 2025), LLMEraser reformulates the calculation of parameter changes as solving an equivalent optimization problem expressed in summation form, enabling efficient resolution using mini-batch algorithms. Specifically, we focus on the following optimization problem regarding $\Delta$:
@@
$f((x,y),\Delta)=\frac{1}{2}\Delta^{\top}\nabla_{\Theta}^{2}\mathcal{L}\left((x,y),\hat{\Theta}\right)\Delta-\langle b,\Delta\rangle.$ (15)
By employing scalable algorithms (e.g., SGD) to optimize problem 12, we can obtain the solution for $\Delta\Theta_{\text{Task}}$. It is worth noting that both the function value and the gradient can be efficiently computed using the Hessian-vector-product (HVP) [^1] HVP has a corresponding implementation in PyTorch; refer to https://pytorch.org/docs/stable/autograd.html for details., reducing the complexity from $\mathcal{O}(p^{2})$ to $\mathcal{O}(p)$ (Pearlmutter, 1994), where $p$ is the number of trainable parameters. The pseudocode for computing parameter changes can be found in Appendix B. Error analysis for our proposed algorithm can be found in Appendix D.
@@ 4.1 Experimental Setups
We conduct experiments on both LLMs and Multimodal Large Language Models (MLLMs), focusing specifically on LLMs for Recommendation (LLM4Rec) (Bao et al., 2023; Liao et al., 2024) and MLLM relation mining tasks (Wu et al., 2024c; Ye et al., 2024), to validate the effectiveness of our proposed LLMEraser. We choose LLaMA2-7B (Touvron et al., 2023b) as our backbone LLM and LLaVA 1.5-7B (Liu et al., 2023a) for the MLLM experiments. Comprehensive details on task, datasets, baselines, and evaluation metrics for our proposed LLMEraser can be found in Appendix C.
@@ 4.4 Results Analysis for Unlearning Efficiency (RQ3)
Efficiency is a key metric in evaluating unlearning techniques, particularly for LLMs. We here conduct experiments, comparing our proposed LLMEraser against existing techniques. For a fair comparison, we report the execution time in the QM task, where $5\%$ of users have items replaced with noisy interactions. All methods are run on a single Nvidia A100 GPU. Table 6 presents the results. We can observe that: (1) Due to the parallel training of sub-models, the retraining time of both SISA and RecEraser can be reduced to some extent. However, RecEraser requires data partitioning based on similarity, which introduces additional computational overhead. Moreover, both methods remain highly inefficient as unlearning requests necessitate retraining of the adapters. (2) In contrast, our proposed LLMEraser exhibits remarkable efficiency in handling unlearning tasks. By directly modifying model parameters, LLMEraser achieves a speedup of approximately 31.25 times compared to retraining, requiring only about $1.4\times 10^{3}$seconds to update the parameters. This reduction in execution time demonstrates the effectiveness of our approach in accelerating the computation of parameter changes. Additional experimental results and related analyses on the memory usage and execution time of LLMEraser can be found in Appendix E.
@@ 6 Conclusion And Future Work
7 Acknowledgements
This research is supported by the National Science and Technology Major Project (2023ZD0121102), National Natural Science Foundation of China (92270114, 62302321, U24B20180, 62121002). The work of Yancheng Yuan is supported by the Research Center for Intelligent Operations Research and The Hong Kong Polytechnic University under grant P0045485. This research is also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
Appendix A Overview of Existing LLM Unlearning Methods
@@
  • APA (Hu et al., 2024c): This study introduces the first exact unlearning approach for large language model-based recommendation (LLMRec), focusing on the removal of personal data to comply with privacy regulations. The Adapter Partition and Aggregation (APA) method is proposed, which combines data partitioning with parameter aggregation to reduce inference latency while maintaining performance. This approach enables efficient unlearning without incurring the extra costs typically associated with traditional methods. However,it can affect the integrity of the adapter structure and necessitates retraining of sub-models.
@@
Appendix B Algorithm for calculating parameter changes
@@
Appendix C Experimental details
@@
  • In terms of MLLMs unlearning tasks, we focus on hard hallucination mining, e.g., understanding of relation (Wu et al., 2024c) and spurious biases (Ye et al., 2024). We structure the evaluation as binary or multi-choice classification problems, which aim to select the ground-truth from the noisy labels. Specifically, for relation understanding, we follow (Wu et al., 2024c) to present the Recall, F1-Score, Precision, Classification accuracy, Yes ratio as the evaluation metrics. For spurious biases, we follow (Wu et al., 2024c) to show the classification accuracy for 9 types of spurious correlations, which is Background (BG), Texture and Noise (TN), Co-occurring Objects (CO), Relative Size (RS), Colorization (Col.), Orientation (Ori.), Lighting and Shadows (LS), Perspective and Angle (PA), and Shape (Sha.).
Datasets: Our experimental datasets for LLM4Rec unlearning tasks include three commonly used recommendation datasets: BookCrossing (Ziegler et al., 2005), MovieLens (Harper & Konstan, 2016), and LastFM (Cantador et al., 2011). We follow the data preprocessing and dataset partitioning as described in (Bao et al., 2023) and (Liao et al., 2024). For MLLMs unlearning tasks, we utilize MMSpuBench (Ye et al., 2024), and R-Bench (Wu et al., 2024c) with the representative masked instances for evaluation, partitioning the data is into training (60%), validation (20%), and testing (20%) set.
@@
Appendix D Estimation Errors Analysis of LLMEraser
The approximation errors in LLMEraser consist of two primary components: first, the errors introduced by the Taylor expansion approximation in the derivation of the influence function, where high-order terms are neglected; and second, the errors arising from the new algorithm proposed in Section 3.3 for solving the inverse Hessian-vector-product. We will conduct the error analysis in two parts accordingly.
D.1 Errors Analysis for Taylor Expansion Approximation
Without loss of generality, we consider approximation error in Equation 6. In other words, we will analyze the error $\|\Delta\Theta(\epsilon)+\epsilon H_{\hat{\Theta}}^{-1}\nabla_{\Theta}\mathcal{L}((x,y);\hat{\Theta})\|.$
The derivation below follows from (Zhang et al., 2022), where we assume that $H_{\hat{\Theta}}$ is invertiable. As we discussed in our paper, this can be guaranteed if the second-order sufficient condition holds at $\hat{\Theta}$.
Since $\hat{\Theta}_{\text{new}}(\epsilon)$ is an optimal solution to the perturbed loss function defined in Equation 5, we have